2018 2017 Accounts receivable $40,000 $36,000 Inventory 28,000 35,000 Net sales 190,000 186,000 Cost of goods...

2018

2017

Accounts receivable

$40,000

$36,000

Inventory

28,000

35,000

Net sales

190,000

186,000

Cost of goods sold

114,000

108,000

Total assets

425,000

405,000

Total stockholders’ equity

240,000

225,000

Net income

32,500

28,000

101) The 2018 debt to equity ratio is (rounded to one decimal place):

A) 80.0%.                         B) 40.0%.                         C) 77.1%.                         D) 60.0%.

Sales revenue

$440,000

Advertising expense

60,000

Interest expense

10,000

Salaries expense

55,000

Utilities expense

25,000

Income tax expense

45,000

Cost of goods sold

180,000

102)

What is operating income?

A) $120,000.                    B) $65,000.                      C) $110,000.                    D) $260,000.

  1. In a perpetual inventory system, the entry at the time of a sale to record the cost of the inventory sold includes a:
    1. Not recorded at the time of the sale.                     B) Debit to Cost of Goods Sold.

C) Credit to Cost of Goods Sold.                              D) Debit to Accounts Receivable.

  1. Goodwill is:
    1. Only recorded by the seller of a business.
    2. Recorded when created internally through advertising expense.
    3. Amortized over the greater of its estimated life or forty years.
    4. The value of a business as a whole, over and above the value of its net identifiable assets.
  1. ABC purchased equipment for $60,000 on January 1, 2018. The equipment is expected to have a five-year life, with a residual value of $5,000 at the end of five years.

Using the straight-line method, the book value at December 31, 2018 would be:

A) $44,000.                      B) $55,000.                      C) $49,000.                      D) $60,000.
Answer #1

1) Debt equity ratio= Total debt/ Total equity

= 185000/ 240000

= 77.1%

So, the answer is option C) 77.1%

2)

Sales revenue 440000
Less: Cost of goods sold (180000)
Gross profit 260000
Less: Operating expenses
Advertising expense (60000)
Salaries expense (55000)
Utilities expense (25000)
Operating income 120000

So, the answer is option A) $120000

3)

Date Account & explanation Debit Credit
Cost of goods sold
Inventory

So, the answer is option B) Debit to Cost of goods sold

4) Goodwill is an intangible asset. It increases the value of the business.

So, the answer is option D) The value of a business as a whole, over and above the value of its net identifiable assets

5) Depreciation using straight-line method= (60000-5000)/ 5= 11000

Book value at December 31, 2018= 60000-11000= 49000

So, the answer is option C) $ 49000